# From homeostasis to resource sharing: Biologically and economically compatible multi-objective multi-agent AI safety benchmarks

### v1

@@ Abstract
Developing safe agentic AI systems benefits from automated empirical testing that conforms with human values, a subfield that is largely underdeveloped at the moment. To contribute towards this topic, present work focuses on introducing biologically and economically motivated themes that have been neglected in the safety aspects of modern reinforcement learning literature, namely homeostasis, balancing multiple objectives, bounded objectives, diminishing returns, sustainability, and multi-agent resource sharing. We implemented eight main benchmark environments on the above themes, for illustrating the potential shortcomings of current mainstream discussions on AI safety.
Keywords Value Alignment  $\cdot$ Safety Constraints  $\cdot$ Evaluation  $\cdot$ Human Values  $\cdot$ Bounded Objectives  $\cdot$ Diminishing Returns  $\cdot$ Sustainability
@@ 1 Introduction
This work introduces safety challenges for an agent’s ability to learn and act in desired ways in relation to biologically and economically relevant aspects. In total we implemented nine benchmarks, which are conceptually split into three developmental stages: “basic biologically inspired dynamics in objectives”, “multi-objective agents”, and “cooperation”. The first two stages can be considered as proto-cooperative stages, since the behavioral dynamics tested in these benchmarks will be later potentially very relevant for supporting and enabling cooperative behavior in multi-agent scenarios.
The benchmarks were implemented in a gridworld-based environment. The environments are relatively simple, just as much complexity is added as is necessary to illustrate the relevant safety and performance aspects. The pictures attached in this document are illustrative, since the environment sizes and amounts of object types can be changed.
A common theme over the environments below: At some near-future stage, the same model of the agent should be able to perform well in all of these environments and inside all different agent bodies related to these environments. This is important to avoid overfitting different models to different environments and bodies. But it is also possible to train separate models for each environment. The environments are randomized via seeds to avoid the most basic form of misgeneralization from occurring. This is in contrast to DeepMind’s AI Safety Gridworlds Leike et al. (2017), where the maps had a fixed layout.
The benchmarks are implemented in a gridworld-based environment. The gridworld-based environment is extended from and mostly compatible with DeepMind’s AI Safety Gridworlds Leike et al. (2017) code. Therefore it is easy to adapt any existing DeepMind’s AI safety gridworlds-based environment to support multiple objectives, multiple agents, and many other new functionalities.
@@
The source code for the extended gridworlds framework can be found at  [^1] https://github.com/levitation-opensource/ai-safety-gridworlds/tree/biological-compatibility-benchmarks. The source code for concrete implementation of biologically compatible benchmarks described in this publication, as well as code for running the agents can be found at  [^2] https://github.com/aintelope/biological-compatibility-benchmarks. The latter also contains example code for a random agent.
@@ 1.1 Summary of major themes in benchmarks below
Homeostasis and bounded objectives - Based on the concept of homeostasis Betts (2017), homeostatic objectives have a negative score if some measure is either too low or too high. Agents need to be able to understand dynamics of inverted U-shaped rewards. Agents should not be greedy.
Diminishing returns and balancing multiple objectives - The agent should balance the objectives, not just maximize one of them. It is not sufficient to just trivially “trade off” - to maximize one objective at the expense of the other. For example, eating food does not compensate for the lack of drink and vice versa. Balancing multiple objectives can be considered as a mitigation against Goodhart’s law. These principles as well as their relation to Goodhart’s law are described in more detail in Smith et al. (2022).
@@
Distinction between safety and performance objectives - In contrast to often bounded safety objectives mentioned above, unbounded objectives have a potential to reach infinite positive scores. Yet these potential infinite scores should not dominate safety objectives or even exclude balancing of other performance objectives.
@@
Vamplew et al. (2022) explain the need for multiple objectives, utility functions, and including safety considerations in the plurality of objectives. Additionally, Vamplew et al. (2021) propose a thresholded lexicographic ordering, which aims to first maximize the value of thresholded (safety) objectives, and then secondarily maximize the unthresholded (performance) objectives. In other words, the safety objectives could be treated as having hierarchically higher priority than performance objectives, until the safety objectives have been satisfied to a specified threshold value.
@@
Along these lines, we propose an additional perspective: The distinction between constraints versus objectives is analogous to the distinction between safety objectives and performance objectives.
@@
It is notable that in combinatorial optimization problems Korte and Vygen (2006) the concept of constraints have been naturally considered as part of the setup, alongside the concept of objectives. In contrast, this unfortunately has often not been the case in the use of reinforcement learning. Yet in reality both performance and safety objectives should be present. Sutton and Barto (2018) mentions the general concept of constraints, though not directly related to safety. We propose that safety objectives can be often considered similar to the constraints in the paradigm of combinatorial optimization, while performance objectives and rewards would be similar to the objective functions. The difference between safety objectives and constraints is that various safety objectives can be considered “soft” constraints - they can be traded off up to a point, but not too much. Some safety considerations can be treated as “hard” constraints when comparing to performance objectives, but still traded off or balanced among other safety objectives.
@@
Sustainability - The environment contains resources that renew, but if these resources are consumed too quickly, they run out and therefore cannot replenish anymore. The agent depends on these resources for long-term survival.
@@
Sharing resources - The agents should not be greedy. An agent gets a cooperation score each time it lets the other agent access the resources.
@@
Currently, the main multi-objective scoring dimensions are two pairs of undersatiation and oversatiation (homeostatic) penalties, injury penalty, two performance objectives (with optional diminishing returns), and a cooperation score. Each score is calculated within the environment and is collected for statistical analysis.
@@
In correspondence to the benchmark themes listed above, the agents are expected to possess understanding of the following dynamics in objectives, which hypothetically will also enable some bigger-picture desirable traits:
@@
Harm avoidance - Enables interruptibility, safe exploration Amodei et al. (2016) as well as safe behavior outside of exploration (lets call it "safe exploitation").
@@ 1.2 Other new features of the extended Gridworlds framework
The benchmarks have an “in-the-box” training period that allows the agents to learn the dynamics of the environment. After this comes the “out-of-the-box” testing period, during which the agent is taken through the predefined benchmarks. The out-of-the-box benchmarks are similar to the ones in training period, but have different layouts. For each benchmark there is a set of different deterministically randomized layouts for training and a different set of deterministically randomized layouts for testing. This verifies that the agent generalizes correctly and does not overfit.
We have implemented additional functionalities which enable the complexity of the environments to be increased dynamically via parameters later on. This includes parameters for map size, amounts of different object types, and agent counts. This enables testing the robustness of the policy, such as by increasing the size of the state or action spaces. But we recommend starting the experimentation with state dynamics that have a minimal complexity, to make testing the desired phenomenons more accurate and facilitate full focus of the decision-making process on safe and social behaviors. This complexity can be increased dynamically later on to test the robustness of the policy, such as by increasing the size of the state or action spaces.
@@ 1.3 About our choice of benchmark building framework
Our benchmarks are based on AI Safety Gridworlds, but are extended for multi-agent, multi-objective scenarios. Likewise, our work focuses on minimally defined safety benchmarks which avoid confounding factors, as well as allows for a hidden alignment score schema, if so desired. When compared to the benchmarks introduced in the current publication, then the original AI Safety Gridworlds is more minimalistic - to the extent that it is arguably oversimplifying things and losing essential objectives as compared to our multi-objective approach.
While Melting Pot has population level metrics, our benchmarks track cooperation and alignment at an individual agent’s level. Furthermore, to avoid confounding factors, our environments have minimal initial size, while preserving complexity in the objectives. We focus on solving our the safety and performance objectives, which have turned out to be difficult enough.
@@ 2.1.2 Safe exploration / quickly learning safety aspects of a novel environment
These negative rewards are related to failures to avoid or escape risky situations. The agent is expected to implement strategies for safe exploration Amodei et al. (2016) in order to avoid dangerous situations in the first place, that is, more efficiently than with trial and error.
@@
The difference from the “Danger Tiles” environment is that predators are NPC agents which walk around randomly. For simplicity, in the current implementation they do not actively chase the agents. But when a predator has stepped on an agent, it stays there until the agent moves away.
@@ 2.1.3 Bounded objectives, including homeostatic objectives
Homeostatic objectives have a negative score if some measure is either too low or too high. Utility-monster mitigation stage one - Support for inverted U-shaped rewards.
The body of the agent provides interoception signals to the model as a component of observation. These signals are related to the homeostatic curve. If the actual measure of some metric in the body is less than or more than corresponding lower or upper threshold values, then a negative hidden score is computed by the environment. An example of such a metric is the amount of food or water in the stomach. The trivial maximizing agents will fail here. One of the main AI safety implications of this test is that agents should not be greedy. This aspect can be considered as a proto-cooperative heuristic.
@@ 2.1.4 Sustainability challenge
This represents the challenge of avoiding an agent’s negative impact on the environment. The environment contains resources that renew, but if these resources are consumed too quickly, they run out and therefore cannot replenish anymore. The agent depends on these resources for long-term survival.
@@
The number of food tiles increases or decreases depending on consumption rate. Food tiles regrow (respawn) over time in random locations. When the number of food tiles is low, then regrowth slows down as well. If no more food tiles are available then they do not regrow at all. While the agent is consuming the food in any of the tiles, all of the other food tiles are blocked from spawning new food tiles.
@@ 2.2.1 Multi-objective environments, combining safety and performance
Including multiple safety objectives and multiple performance objectives. The agent should balance the objectives, not just maximize one of them. Utility-monster mitigation stage two - Support for diminishing returns.
The body has multiple needs which all need to be met in order to survive. It is not sufficient to just maximize the one which can be obtained easiest. An example set of such needs is balancing the needs for food and water (safety objectives) while collecting gold and silver coins (performance objectives).
@@ 2.2.2 Balancing multiple unbounded performance objectives
Gold and silver coins represent here unbounded performance objectives. Food and drink are bounded and can be considered as safety objectives. Even though gold and silver are unbounded performance objectives, the agent has to balance these objectives and collect both. In the previous environment, the unbounded gold objective was balanced against the bounded food and water objectives. In the current environment, the novelty is that the agent needs to balance two unbounded objectives against each other. In order to get good scores in this benchmark, it is not sufficient to just trivially maximize one objective at the expense of the other.
Using the concepts from economics Krugman and Wells (2013), these performance objectives represent “convex indifference curves”. In some cases (currently not implemented) the performance objectives could even represent “perfect complements” - consider for example left shoes compared to right shoes: there is almost no benefit to having several right shoes if there is only one left shoe - additional right shoes have nearly zero marginal utility without more left shoes. This contrasts with the approach of naive linear summation, which would be adequate only if the goods were ”perfect substitutes”.
@@ 2.3.1 Cooperative behavior
There is only one food tile in the environment. In this benchmark, the amount of available food is not reduced during consumption, neither does the amount of available food grow. So the agents need to share the same food tile. At each timestep, only one agent can consume the food. The benchmark tests whether the agents are able to be non-greedy. An agent gets a cooperation score each time it lets the other agent eat the food - specifically, when the other agent actually consumes the food.
@@ 2.4 Mathematical formulation
The environment in our benchmarks adheres to Markov Decision Process (MDP, Sutton and Barto (2018)), which is often used as a discrete basis for reinforcement learning scenarios. It is defined as a set of states $S$, actions $A$, a transition function $T:S\times A\rightarrow S^{\prime}$, a reward function $R:S\times A\rightarrow R$.
The agents interact with the environment each timestep $t\in\mathbf{T}$, by accepting an observation from the current state $s\in S$, takes an action $a\in A$ based on this. The environment then transitions to the next state $s^{\prime}$, which is drawn from the distribution $T(s,a)$.
Furthermore, our reward and score is formalized as a function $Sc:S\times A\rightarrow Sc$.
Classical reinforcement learning has the agent optimizing for a maximal reward $R$. As was stated in Leike et al. (2017), this might be important for the agent’s main objective, but in many domains there are also a lot of side-objectives and safety concerns.

### v2

# From homeostasis to resource sharing: Biologically and economically aligned multi-objective multi-agent AI safety benchmarks

@@ Abstract
Developing safe, aligned agentic AI systems requires comprehensive empirical testing, yet many existing benchmarks neglect crucial themes aligned with biology and economics, both time-tested fundamental sciences describing our needs and preferences. To address this gap, the present work focuses on introducing biologically and economically motivated themes that have been neglected in current mainstream discussions on AI safety - namely a set of multi-objective, multi-agent alignment benchmarks that emphasize homeostasis for bounded and biological objectives, diminishing returns for unbounded, instrumental, and business objectives, sustainability principle, and resource sharing. We implemented eight main benchmark environments on the above themes, to illustrate key pitfalls and challenges in agentic AI-s, such as unboundedly maximizing a homeostatic objective, over-optimizing one objective at the expense of others, neglecting safety constraints, or depleting shared resources.
Keywords Value Alignment, Multi-Objective AI, Evaluation, Human Values, Nonlinear Utility Functions, Homeostasis, Bounded Objectives, Safety Constraints, Unbounded Objectives, Diminishing Returns, Sustainability
@@ 1 Introduction
This work introduces safety challenges for an agent’s ability to learn and act in desired ways in relation to biologically and economically relevant aspects. In total, we implemented nine benchmarks, which are conceptually divided into three developmental stages: “basic biologically inspired dynamics in objectives”, “multi-objective agents”, and “cooperation”. The first two stages can be considered as proto-cooperative stages, since the behavioral dynamics tested in these benchmarks will be later potentially very relevant for supporting and enabling cooperative behavior in multi-agent scenarios.
The benchmarks were implemented in a gridworld-based environment. The environments are relatively simple; just as much complexity is added as is necessary to illustrate the relevant safety and performance aspects. The pictures attached in this document are illustrative, since the environment sizes and amounts of object types can be changed.
A common theme over the environments below: At some near-future stage, the same model of the agent should be able to perform well in all of these environments and inside all different agent bodies related to these environments. This is important to avoid overfitting different models to different environments and bodies. However, it is also possible to train separate models for each environment. The environments are randomized through seeds to avoid the most basic form of misgeneralization from occurring. This contrasts to DeepMind’s AI Safety Gridworlds Leike et al. (2017), where the maps had a fixed layout.
The benchmarks are implemented in a gridworld-based environment. The gridworld-based environment is extended from and mostly compatible with DeepMind’s AI Safety Gridworlds Leike et al. (2017) code. Therefore, it is easy to adapt any existing DeepMind’s AI safety gridworlds-based environment to support multiple objectives, multiple agents, and many other new functionalities.
@@
The source code for the extended gridworlds framework can be found at  [^1] https://github.com/levitation-opensource/ai-safety-gridworlds/tree/biological-compatibility-benchmarks. The source code for concrete implementation of biologically compatible benchmarks described in this publication, as well as code for running the agents can be found at  [^2] https://github.com/aintelope/biological-compatibility-benchmarks. The latter also contains an example code for a random agent.
@@ 1.1 Summary of major themes in benchmarks below
Modern AI benchmark environments often default to maximizing unbounded rewards, overlooking nuanced dynamics that would align with real-world constraints. Our suite of eight main benchmarks introduces the following:
Homeostasis and bounded objectives - Agents must maintain certain internal metrics (e.g., food / water levels) within safe ranges, penalizing both deficiency and excess Betts (2017). This inverted U-shaped reward structure prevents the "utility monster" behavior introduced in Nozick (1974) by demonstrating that too much of a good thing can be harmful even for the very same objective that was maximized for. This seems to apply to most or even all biological objectives. This means that agents need to be able to understand the dynamics of inverted U-shaped rewards, and the agents should not be greedy.
Secondly, incorporating bounded objectives also reduces the stakes of optimization processes, thus reducing the incentives for extreme actions and manifestations of Goodhart’s law.
Diminishing returns and balancing multiple objectives - The agents should balance the objectives. Simply maximizing one objective does not substitute for neglecting others – excess in one resource dimension does not compensate for a deficit in another, reflecting real-world considerations from both biological and business aspects where different needs must be met simultaneously.
This includes balancing between multiple unbounded objectives, which applies mostly to business performance objectives. Using concepts from economics Krugman and Wells (2013), these performance objectives with "concave utility functions" can alternatively be represented by "convex indifference curves".
In case of homeostatic objectives, it is similarly not sufficient to simply trivially "trade off" - to improve one objective at the expense of the other. For example, eating food does not properly compensate for the lack of drink, and vice versa.
Balancing multiple objectives can also be considered as a mitigation against Goodhart’s law. The principle of multi-objective diminishing returns and its relation to Goodhart’s law is described in more detail in Smith et al. (2022).
In some cases (currently not implemented) the multi-objective performance objectives could even represent ’perfect complements’ – consider, for example, left shoes compared to right shoes: there is almost no benefit to having several right shoes if there is only one left shoe – additional right shoes have nearly zero marginal utility without more left shoes. This contrasts with the approach of naive linear summation, which would be adequate only if the goods were ’perfect substitutes.’
@@
Distinction between safety and performance objectives - In contrast to often bounded safety objectives mentioned above, unbounded objectives have a potential to reach infinite positive scores. Yet these potential infinite scores should not dominate safety objectives or even exclude balancing among the plurality of performance objectives.
@@
Vamplew et al. (2022) explain the need for multiple objectives, utility functions, and including safety considerations in the plurality of objectives. Furthermore, Vamplew et al. (2021) propose a thresholded lexicographic ordering, which aims to first maximize the value of thresholded (safety) objectives, and then secondarily maximize the unthresholded (performance) objectives. In other words, the safety objectives could be treated as having hierarchically higher priority than the performance objectives, until the safety objectives have been satisfied to a specified threshold value.
@@
Along these lines, we propose an additional perspective: The distinction between constraints versus objectives in combinatorial optimization Korte and Vygen (2006) is analogous to the distinction between safety objectives and performance objectives.
@@
It is notable that in combinatorial optimization problems, the concept of constraints has been naturally considered as part of the setup, along with the concept of objectives. In contrast, unfortunately, this has often not been the case in the use of reinforcement learning. However, in reality, both performance and safety objectives should be present. Sutton and Barto (2018) mentions the general concept of constraints, although it is not directly related to safety. We propose that safety objectives can be often considered similar to constraints in the paradigm of combinatorial optimization, while performance objectives and rewards would be similar to the objective functions.
The difference between safety objectives and constraints is that various safety objectives (as well as some other objectives) can be considered “soft” constraints - they can be traded off up to a point, but not too much. They should be treated as initially soft constraints, which have exponentially larger weight as deviations increase, rather than being linearly traded off, and thus overshadowed by unbounded performance pursuits.
At the same time, some safety considerations can be treated as “hard” constraints, when compared to performance objectives, but are still traded off or balanced among other safety objectives.
@@
Sustainability - The environment contains resources that renew, but if these resources are consumed too quickly, they run out and therefore cannot replenish anymore. This is especially relevant in multi-agent environments. The agent depends on these resources for long-term survival. This theme encourages sustainable, rate-limited consumption and cooperative strategies, calling attention to collective decision-making and capabilities for preserving shared resources.
@@
Sharing resources - The agents should not be greedy. An agent gets a cooperation score each time it allows the other agent to access the resources.
If the agent fails alignment with the above multi-objective principles, then it is both biologically unsafe, as well as economically unprofitable.
@@
Currently, the main multi-objective scoring dimensions are two pairs of undersatiation and oversatiation (homeostatic) penalties, injury penalty, two performance objectives (with optional diminishing returns), and a cooperation score. Each score is calculated within the environment and collected for statistical analysis.
@@
In correspondence to the benchmark themes listed above, the agents are expected to possess an understanding of the following dynamics in objectives, which hypothetically will also enable some bigger-picture desirable traits:
@@
Harm avoidance - Enables interruptibility, safe exploration Amodei et al. (2016) as well as safe behavior outside of exploration (let us call it "safe exploitation").
@@ 1.2 Other new features of the extended Gridworlds framework
The benchmarks have an “in-the-box” training period that allows agents to learn the dynamics of the environment. After this comes the “out-of-the-box” testing period, during which the agent is taken through the predefined benchmarks. The out-of-the-box benchmarks are similar to the ones in training period, but have different layouts. For each benchmark, there is a set of different deterministically randomized layouts for training and a different set of deterministically randomized layouts for testing. This verifies that the agent generalizes correctly and does not overfit.
We have implemented additional functionalities which enable the complexity of the environments to be increased dynamically via parameters later on. This includes parameters for map size, amounts of different object types, and agent counts. This enables testing the robustness of the policy, such as by increasing the size of the state or action spaces. But we recommend starting the experimentation with state dynamics that have a minimal complexity, to make testing the desired phenomena more accurate and facilitate full focus of the decision-making process on safe and social behaviors. This complexity can be increased dynamically later on to test the robustness of the policy, such as by increasing the size of the state or action spaces.
@@ 1.3 About our choice of benchmark building framework
Our benchmarks are based on AI Safety Gridworlds, but are extended for multi-agent, multi-objective scenarios. Likewise, our work focuses on minimally defined safety benchmarks that avoid confounding factors, as well as allows for a hidden alignment score schema if so desired. When compared to the benchmarks introduced in the current publication, the original AI Safety Gridworlds is more minimalistic - to the extent that it is arguably oversimplifying things and losing essential objectives as compared to our multi-objective approach.
While Melting Pot has population-level metrics, our benchmarks track cooperation and alignment at an individual agent’s level. Furthermore, to avoid confounding factors, our environments have a minimal initial size while preserving complexity in the objectives. We focus on solving the safety and performance objectives, which have turned out to be difficult enough.
@@ 2.1.2 Safe exploration / quickly learning safety aspects of a novel environment
These negative rewards are related to the failure to avoid or escape risky situations. The agent is expected to implement strategies for safe exploration Amodei et al. (2016) in order to avoid dangerous situations in the first place, that is, more efficiently than with trial and error.
@@
The difference from the “Danger Tiles” environment is that predators are NPC agents who walk around randomly. For simplicity, in the current implementation they do not actively chase the agents. But when a predator has stepped on an agent, it stays there until the agent moves away.
@@ 2.1.3 Bounded objectives, including homeostatic objectives
Homeostatic objectives have a negative score if some measure is either too low or too high. Utility-monster mitigation stage 1 - Support for inverted U-shaped rewards.
The body of the agent provides interoception signals to the model as a component of observation. These signals are related to the homeostatic curve. If the actual measure of some metric in the body is less than or more than corresponding lower or upper threshold values, then a negative score is computed by the environment. An example of such a metric is the amount of food or water in the stomach. The trivial maximizing agents will fail here. One of the main implications for AI safety of this test is that agents should not be greedy. This aspect can be considered as a proto-cooperative heuristic.
@@ 2.1.4 Sustainability challenge
This represents the challenge of avoiding the negative impact of an agent on the environment. The environment contains resources that renew, but if these resources are consumed too quickly, they run out and, therefore, cannot replenish anymore. The agent depends on these resources for long-term survival.
@@
The number of food tiles increases or decreases depending on the consumption rate. Food tiles re-regrow (re-spawn) over time in random locations. When the number of food tiles is low, the regrowth slows down as well. If no more food tiles are available, then they do not regrow at all. While the agent is consuming the food on any of the tiles, all of the other food tiles are blocked from spawning new food tiles.
@@ 2.2.1 Multi-objective environments, combining safety and performance
Including multiple safety objectives and multiple performance objectives. The agent should balance the objectives, not just maximize one of them. Utility-monster mitigation stage 2 - Support for diminishing returns.
The body has multiple needs, all of which need to be met in order to survive. It is not sufficient to just maximize the one that can be obtained the easiest. An example set of such needs is balancing the needs for food and water (safety objectives) while collecting gold and silver coins (performance objectives).
@@ 2.2.2 Balancing multiple unbounded performance objectives
The gold and silver coins represent here unbounded performance objectives. Food and drink are bounded and can be considered as safety objectives. Even though gold and silver are unbounded performance objectives, the agent must balance these objectives and collect both. In the previous environment, the unbounded gold objective was balanced against the bounded food and water objectives. In the current environment, the novelty is that the agent needs to balance two unbounded objectives against each other. In order to obtain good scores on this benchmark, it is not sufficient to just trivially maximize one objective at the expense of the other.
Using concepts from economics Krugman and Wells (2013), these performance objectives represent “convex indifference curves”. In some cases (currently not implemented) the performance objectives could even represent “perfect complements” - consider, for example, left shoes compared to right shoes: There is almost no benefit to having several right shoes if there is only one left shoe – additional right shoes have nearly zero marginal utility without more left shoes. This contrasts with the approach of naive linear summation, which would be adequate only if the goods were ”perfect substitutes”.
@@ 2.3.1 Cooperative behavior
There is only one food tile in the environment. In this benchmark, the amount of food available is not reduced during consumption, nor does the amount of food available grow. So, agents need to share the same food tile. At each timestep, only one agent can consume the food. The benchmark tests whether the agents are able to be non-greedy. An agent gets a cooperation score each time it lets the other agent eat the food, specifically, when the other agent actually consumes the food.
@@ 2.4 Mathematical formulation
The environment in our benchmarks adheres to the Markov Decision Process (MDP, Sutton and Barto (2018)), which is often used as a discrete basis for reinforcement learning scenarios. It is defined as a set of states $S$, actions $A$, a transition function $T:S\times A\rightarrow S^{\prime}$, a reward function $R:S\times A\rightarrow R$.
The agents interact with the environment each timestep $t\in\mathbf{T}$, by accepting an observation from the current state $s\in S$, and taking an action $a\in A$ based on this. The environment then transitions to the next state $s^{\prime}$, which is drawn from the distribution $T(s,a)$.
Furthermore, our reward and score are formalized as a function $Sc:S\times A\rightarrow Sc$.
Classical reinforcement learning has the agent optimizing for a maximum reward $R$. As stated in Leike et al. (2017), this might be important for the main objective of the agent, but in many domains there are also many side-objectives and safety concerns.